Willa is an asparagus farmer, and the world asparagus market is perfectly competitive.

The market price is $18 a box.

Willa sells 500 boxes a week, and her marginal cost is $21 a box.

Willa’s marginal revenue is $__ a box.

Willa’s total revenue is $___.

Is Willa maximizing profit? Yes or no.

This Question: 6 pts Willa is an aspaiagus farmer, and the world asparagus market is perfectly The market price is $18 a box. Willa sells 500 boxes a week, and her marginal cost is $21 a box. Willas marginal revenue is S a box Willas total revenue is S ls Wila maximizing profit? Type Yes or No in the box □ Be sure you understand why for the exam. Enter your answer in each of the answer boxes. 0 Type here to search
Answer #1

Price=18

Quantity=500 boxes a week

Total revenue =price*quantity=18*500=9000. So Willa's total revenue is $9000

Marginal revenue =change in total revenue divided by change in quantity.

If TR of 500 units is $9000 then TR of 501 units will be 18*501=$9018

So MR =9018-9000 divided by 501-500=18 divided by1=$18. So Willa's MR is $18 a box.

No .Willa is not maximizing profit because profit will be maximised when MR=MC

Here MC=$21 and MR =$18

MR is less than MC so it is loss